OpenStudy (anonymous):

A freight train ans a passenger train start toward each other at the same time from two cities that are 600 mile apart. After 4 hours the trains are still 80 miles apart. If the average speed of the freight train is 20 miles per hour faster then the average rate of speed of the passenger train, what is the average rate of speed, in miles per hour, of the passenger train?

OpenStudy (anonymous):

D1 + D2 = 600 - 80 = 520 mi

OpenStudy (anonymous):

Let s is speed of passenger train then s + 20 is speed of freight train @ArkGoLucky Can you build up the equation from here?

OpenStudy (anonymous):

4s+4(s+20)=520

OpenStudy (anonymous):

s=55

OpenStudy (anonymous):

You can double check yourself by plugging back to the original d1 + d2 = 520 miles

OpenStudy (anonymous):

okay thanks